Caroline has 16 coins, all sickles and galleons. Each sickle is worth 29 knuts, and each galleon is work 493 knuts. The total value of her money is 2784 knuts. How many of each type of coins does she have?
Answer by edjones(8007)

Let s=number of sickles and g=number of galleons.
29s+493g=2784
s+g=16
s=16-g
29(16-g)+493g=2784
464-29g+493g=2784
464+464g=2784
464g=2320
g=5
s=11
.
Ed